Lingenfelter Corporation experienced a fire on December 31, 2017, in which its financial records were partially destroyed. It has been able to salvage some of the records and has ascertained the following balances. December 31, 2017 December 31, 2016 Cash $ 27,000 $ 14,500 Accounts receivable (net) 74,500 125,000 Inventory 199,000 185,000 Accounts payable 47,000 91,000 Notes payable 27,500 57,000 Common stock, $100 par 399,000 399,000 Retained earnings 118,000 107,000

Additional information: 1. The inventory turnover is 3.8 times.

2. The return on common stockholders’ equity is 26%. The company had no additional paid-in capital.

3. The accounts receivable turnover is 8.7 times.

4. The return on assets is 12.5%.

5. Total assets at December 31, 2016, were $593,000.

Compute the following for Lingenfelter Corporation. (Round answers to 0 decimal places, e.g. 5,275.)

(a) Cost of goods sold for 2017

(b) Net sales (credit) for 2017

(c) Net income for 2017 (

d) Total assets at December 31, 2017
Answer #1

a) Average inventory= ($199000+185000)/2= $192000

Inventory turnover ratio= Cost of goods sold/Average inventory

3.8= Cost of goods sold/192000

Cost of goods sold= 192000*3.8

Cost of goods sold= $729600

b) Average accounts receivable= ($74500+125000)/2= $99750

Accounts receivable turnover= Net sales/Average accounts receivable

8.7= Net sales/99750

Net sales= $99750*8.7

Net sales= $867825

c) Average common stockholders' equity= ($399000+399000+118000+107000)/2= $511500

Return on common stockholders' equity= Net income/Average common stockholders' equity

26%= Net income/511500

Net income= $511500*26%

Net income= $132990

d) Return on assets= Net income/Average total assets

12.5%= $132990/Average total assets

Average total assets= $132990/12.5%

Average total assets= $1063920

Average total assets= (Beginning assets+Ending assets)/2

$1063920= ($593000+Ending assets)/2

$2127840= $593000+Ending assets

Ending assets= 2127840-593000= $1534840

Total assets at December 31, 2017= $1534840

NOTE:- For any problem regarding the answer please ask in the comment section.